OpenStudy (anonymous):

If (w+3)(2w-a) = bw^2 - w + 3c, then the value of a-b+c will be A) 7 B) -7 C) 8 D) -9 E) -2 'w' maybe OMEGA, but i am not sure.. Help!

OpenStudy (anonymous):

\[\left( w+3 \right)\left( 2w-a \right)=bw^2-w+3c~or~2w^2+6w-aw-3a=bw^2-w+3c\] 2=b -1=6-a -3a=3c from these equations find the values of a,b,c.

OpenStudy (anonymous):

thanks...

OpenStudy (anonymous):

yw